OpenStudy (anonymous):

HELP!! Which statement is true? A.–16 + 23 = –17 + 10 B.15 + (–5) > 23 + (–12) C.8 + (–7) < –15 + 14 D.–12 + (–11) > 5 + (–29)

OpenStudy (anonymous):

Okay first off simplify

OpenStudy (anonymous):

-16+23=? In other words 23-16=7 Do the same for the second part of the problem -17+10=-7 First one doesn't work... Now do the rest on your own...

OpenStudy (anonymous):

Its pretty easy

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

15-5=10 23-12=11 second one doesn't work either

OpenStudy (anonymous):

8-7=1 -15+14=1 Doesn't work 1 is not greater than 1

OpenStudy (anonymous):

Now since you only have 1 answer left that it could be what do you think it is?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

-23>-24

OpenStudy (anonymous):

D Correct

OpenStudy (anonymous):

Because owing someone 23 dollars is better than owing them 24...